OpenStudy (anonymous):

10square root 125 simplify

OpenStudy (solomonzelman):

\[\huge\color{blue}{ 10 \sqrt{125} } \] \[\huge\color{red}{ HINT:~~~~~125=5^3=5 \times 5 \times 5 } \]

OpenStudy (anonymous):

So the answer would be 3 square root 5? Sorry to ask such dumb questions I'm not very good at this it's all my first time learning it

OpenStudy (solomonzelman):

I'll show you on a different example, here,\[\huge\color{blue}{ 6 \sqrt{27} } \]\[\huge\color{blue}{ 6 \sqrt{3^3} } \]\[\huge\color{blue}{ 6 \sqrt{3^2 \times 3} } \]\[\huge\color{blue}{ 6 \sqrt{3^2} \times \sqrt{3} } \]\[\huge\color{blue}{ 6 \times 3 \times \sqrt{3} } \]\[\huge\color{blue}{ 18 \times \sqrt{3} } \]\[\huge\color{blue}{ 18 \sqrt{3} } \]

OpenStudy (solomonzelman):

And use the hint I gave you above, in RED.

OpenStudy (cggurumanjunath):

50*SQRT (3) IS THE FINAL ANSWER.

OpenStudy (solomonzelman):

@CGGURUMANJUNATH why are you giving away the answer, doing this without any explanations?

OpenStudy (anonymous):

I don't get what your doing with the colors?